# A Portable and Stealthy Inaudible Voice Attack Based on Acoustic Metamaterials

### v1

@@ Abstract
We present MetaAttack, the first approach to leverage acoustic metamaterials for inaudible attacks for voice control systems. Compared to the state-of-the-art inaudible attacks requiring complex and large speaker setups, MetaAttack achieves a longer attacking range and higher accuracy using a compact, portable device small enough to be put into a carry bag. These improvements in portability and stealth have led to the practical applicability of inaudible attacks and their adaptation to a wider range of scenarios. We demonstrate how the recent advancement in metamaterials can be utilized to design a voice attack system with carefully selected implementation parameters and commercial off-the-shelf components. We showcase that MetaAttack can be used to launch inaudible attacks for representative voice-controlled personal assistants, including Siri, Alexa, Google Assistant, XiaoAI, and Xiaoyi. The average word accuracy of all assistants is 76%, with a range of 8.85 m.
Index Terms:
Inaudible Attack, Acoustic Metamaterials, Voice Control Systems, Security and Privacy.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I INTRODUCTION
Voice control is a natural way for human-computer interactions, commonly found in mobile and smart home applications [1, 2, 3, 4, 5, 6]. Given its capacity to control a broad spectrum of tasks – from making phone calls to managing devices like surveillance cameras and even initializing bank transfers – voice control systems have increasingly become the target of attacks [7, 8, 9, 10, 11, 12, 13, 6].
Recent research has demonstrated that the voice control mechanisms in mainstream personal assistant systems, like Amazon’s Alexa and Apple’s Siri, are susceptible to inaudible attacks [7, 8, 9]. Such attacks modulate voice commands onto ultrasonic carriers, generating commands imperceptible to human ears but can be accepted by voice control systems. Despite some levels of success in specific scenarios, existing attacking methods have a short attack range of up to 3 m. However, a close-in attack within 3 m can be easily detected by many vigilant individuals [14]. Extending this range by simply increasing transmission power would lead to audible noise, thereby compromising the stealthiness of the attack. More recent work, as represented by LipRead [14], divides the spectrum of the inaudible commands to reduce audible leakage and utilizes loudspeaker arrays to expand the attack range. LipRead is shown to be able to launch inaudible attacks from distances up to 7.62 m using 61 loudspeakers. Another representative system, DolphinAttack [15], can reach 19.8 m for one model of the device (iPhone X), requires 40 speakers, and uses an external power supply and amplification equipment. However, the division of the spectrum of inaudible commands reduces the accuracy of inaudible commands, and the size of the amplification equipment makes it difficult to launch the attack in many practical public scenarios.
@@
Figure 1: A possible attacking scenario of MetaAttack where an attacker can launch an inaudible attack in a public place to remotely control the victim device in another building. In this scenario, the attack device must be stealthy and does not require the use of other continuously powered devices such as DC power supply.
@@
We showcased that MetaAttack can be implemented with low-cost, commercial, off-the-shelf components like Raspberry Pi, amplifiers, loudspeakers, servo, and a 3D-printed MAM. We thoroughly evaluate MetaAttack under different practical scenarios under different attack ranges, tasks, noise levels, obstacles, and targeting different victim devices. We compare MetaAttack against three prior inaudible attacks [8, 9, 14]. Experimental results show that MetaAttack is robust under different evaluation scenarios and tasks and can successfully launch an audible attack within a 8.85 m range with a noise level of up to 65 dB. Compared with prior attacking methods, MetaAttack achieves the good accuracy and attack range (76% average accuracy under a 8.85 m attack range) with the smallest sized device.
@@ II-A Inaudible Attacks
Specifically, when microphones used in voice control systems receive input signals exceeding 20 kHz, due to the non-linearity effect, the microphones output additional non-linear signals ${S_{in}}^{2}(t)$ [15]. ${S_{in}}(t)$ as shown in Eq.1:
@@
where ${V_{attack}}\left(t\right)$ is the voice control signal, and $\cos\left({2\pi{f_{carrier}}t}\right)$ is the high frequency carrier signal. When the attack signal is input to the microphones, the microphones output the signal $S_{attack}(t)$ as shown in Eq.2:
@@
where ${V_{attack}}\left(t\right)$ can be recognized by the microphones, allowing the inaudible attacks to be successfully realized.
@@
The effectiveness of inaudible attacks is compromised by the inherent limitation of a limited attack range compared to laser [34] and wireless signal attacks [35], typically ten feet [8, 9]. This limitation results from the audible leakage at the point of transmission during inaudible attacks. Concealment requires a reduction in output power [14], thereby limiting the range of inaudible attacks. As a result, the impact of an inaudible attack decreases significantly as the distance between the sonic emitter and the target increases, making it less suitable for long-range targeting than other forms of attack.
@@ II-B Threat Model
We focus on implementing an inaudible acoustic attack that is portable, miniaturized, and out of detection range. These characteristics address the need for attacks in portable and stealthy environments. Our attack does not require the attacker to access the victim’s device physically. However, we assume the attacker can synthesize the target’s voice using, e.g., deep-learning-based voice generation from a few training samples [8]. We believe this assumption is reasonable in many day-to-day scenarios where the victim’s voice samples can be collected from public places, a multi-user meeting, or audio and video clips available to the attacker. We also assume that the victim was not intact with the device during the attack and surroundings are not completely enclosed so that it is possible to transmit ultrasounds.
@@ II-C Practical Limitations of Prior Methods
Launching inaudible voice commands involves modulating the control voice signal. However, nonlinear effects from devices like loudspeakers can cause audible leakage, particularly when operating at high power levels, to extend the attack range. Previous works [8, 9] limited device power to prevent sound leakage, but doing so limits the attacking range to around 3 m. To enable longer-range inaudible attacks, LipRead [14] segments the inaudible command spectrum into low-frequency parts that are less detectable by humans. While representing an improvement, LipRead also has drawbacks. It requires an array of 61 loudspeakers to reach up to an attacking distance of 7.62 m, increasing the risk of detection, and the method of segmenting the spectrum may lower command accuracy due to potential recombination or interference during transmission. Similarly, although Yan et al. [15] were able to extend the ultrasonic attack range to 19.8 m, the attack needed a signal generator and a constant DC power supply, which makes it less practical for portable mobile attacks. Additionally, the use of 40 loudspeakers was necessary for the attack device. To more clearly illustrate the limitations of existing studies regarding the distance of the inaudible sound attack, we present a visual representation in Table I. Most of the previous research has been on defeating voiceprint recognition [14, 7, 9, 8], but these studies are difficult to conduct in real-world scenarios. Some of the work requires close proximity to the attacking device, and although long-range attacks can be achieved, they require the introduction of more auxiliary devices [15, 14, 7] that are not stealthy. A viable attack therefore requires a balance between lightness and range.
@@
System name System setup DC power supply Signal generator Speaker number Attack range Portability Stealthy
DolphinAttack (Long-range System) [15] [Uncaptioned image] Yes (Approximately 11 MacBook Pro) Yes (Approximately 12 MacBook Pro) 40 19.8m (iPhoneX) 10m (iPhoneSE) No No
LipRead [14] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 61 7.62m No No
BackDoor [7] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 18 3.5m No No
EchoAttack [9] [Uncaptioned image] No Yes (Approximately 7 MacBook Pro) 40 2.23m Yes Yes
@@ III OVERVIEW of OUR APPROACH
Figure 3: Overview of MetaAttack with three steps: Processing inaudible commands, Attacking victim’s device, and Activating feedback mechanism.
@@
Processing inaudible commands. The initial phase of our attack is to process and modulate voice commands. In this work, we first captured the target’s voice and synthesize the target’s voiceprint (see Section II-B). Then we modulated the voice command groups generated by Google Text-to-Speech AI [39] and voiceprints onto ultrasound carrier waves using the modulation method discussed in Section II to form inaudible commands [14]. Each voice command group used for the attack contains three types of inaudible instructions: commands for muting the victim devices, commands for executing attacks, and commands for feedback. For instance, mute commands might include “speak 6% ($MuteCmd$)” for Siri or “turn the volume to 1 ($MuteCmd$)” for other devices [35, 40], attack commands might include “send message to person X ($AttackCmd$),” while feedback commands could be “turn on hotspot ($FeedbackCmd_{1}$)” and “turn off hotspot ($FeedbackCmd_{2}$)”. Subsequently, a group of inaudible commands is transmitted through a loudspeaker array. During this process, any audible leakage is filtered out, and our MAM amplifies the inaudible commands.
Attacking victim’s device. MetaAttack uses beamforming to send the inaudible commands, but this requires the signal to be transmitted within a specific angle of approximately 12°. To position the signal transmission angle, we use a servo to drive MetaAttack to point to the victim device within a 180° range in steps of 12°. We first send $MuteCmd$ to lower the victim device’s loudspeakers to an inaudible level[35, 40]. Then, we send $AttackCmd$ to achieve the purpose of the attack. Finally, we send $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to activate the feedback mechanism.
Activating feedback mechanism. The feedback mechanism uses the Raspberry Pi to detect the state changes caused by the wireless signal switching command received by the target device, thereby evaluating whether the attack is successful. If the hotspot information of the device completes state transitions between appearance and disappearance in the detection list of MetaAttack after the attacker sends a command group, we consider the execution of the command group successful. Considering that the SSID list within the attack range is not static, the SSID information from weak wireless signals of non-target devices may randomly appear or disappear, we use iwlist [41] to monitor the SSID and signal strength of wireless signals in real-time. Specifically, when the SSID information from weak wireless signals appears or disappears, its signal strength typically shows a gradual increase or decrease. In contrast, the wireless signal from the victim device triggered by $FeedbackCmd$ does not undergo such gradual changes when it appears or disappears. Therefore, after sending $FeedbackCmd$, we use iwlist to identify signals that appear or disappear directly, thereby eliminating the interference from weak wireless signals in our assessment.
@@
As shown in Algorithm 1 and Fig. 4, we determine the attack’s success by identifying the hotspots of smartphones or the Bluetooth connectivity of smart home devices. To illustrate the feedback mechanism for smartphones, we break it down into four sequential steps:
Step 1. After completing the transmission of the $MuteCmd$ and $AttackCmd$ (line 1), the Raspberry Pi uses iwlist to record the identified hotspot devices and stores the information in the L1 list (line 1). The system then proceeds to send the $FeedbackCmd_{1}$ to store the records in the L2 list (line  1).
@@
Step 2. The system continues to send $FeedbackCmd_{2}$ and stores the received hotspot information in list L3 (line  1). If there are discrepancies in the hotspot signals ($dif_{1}$) between the L2 and L3 lists, the process moves to Step 3 (lines  11). Otherwise, the MetaAttack is rotated to the next angle (line  1).
@@
Step 3. MetaAttack will sequentially send $FeedbackCmd_{1}$ and then instruct the Raspberry Pi to retrieve the L4 hotspot list (line  1). If there are differences in the hotspot signals ($dif_{2}$) between the L3 and L4 lists, and the intersection of $dif_{1}$ and $dif_{2}$ contains only one device, a successful stealth attack is detected (lines  11). If not, the MetaAttack will be adjusted to the next angle (line  1).
@@
Step 4. To maintain the stealth of the attack, this step focuses on restoring the victim to its original state before the attack (lines  11).
@@
After these four steps, the system’s attack feedback mechanism ensures the success rate of the inaudible attack. To confirm the feasibility of our attack process, we have released a complete process demonstration [31] and open-sourced the feedback code [33].
Figure 4: Feedback overview of MetaAttack (After using $MuteCmd$ and $AttackCmd$ to achieve the purpose of inaudible attacks, using $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to provide feedback)
Input: $AttackCmd,\ FeedbackCmd_{1},\ FeedbackCmd_{2}$
[…]
2       $Send\ MuteCmd\ and\ AttackCmd$
3       $Activate\ iwlist$

### v2

@@ Abstract
We present MetaAttack, the first approach to leverage acoustic metamaterials for inaudible attacks for voice control systems. Compared to the state-of-the-art inaudible attacks requiring complex and large speaker setups, MetaAttack achieves a longer attacking range and higher accuracy using a compact, portable device small enough to be put into a carry bag. These improvements in portability and stealth have led to the practical applicability of inaudible attacks and their adaptation to a wider range of scenarios. We demonstrate how the recent advancement in metamaterials can be utilized to design a voice attack system with carefully selected implementation parameters and commercial off-the-shelf components. We showcase that MetaAttack can be used to launch inaudible attacks for representative voice-controlled personal assistants, including Siri, Alexa, Google Assistant, XiaoAI, and Xiaoyi. The average success rate of all assistants is 76%, with a range of 8.85 m.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I INTRODUCTION
Voice control is a natural way for human-computer interactions, commonly found in mobile and smart home applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Given its capacity to control a broad spectrum of tasks – from making phone calls to managing devices like surveillance cameras and even initializing bank transfers – voice control systems have increasingly become the target of attacks [11, 12, 13, 14, 15, 16, 17, 1].
Recent research has demonstrated that the voice control mechanisms in mainstream personal assistant systems, like Amazon’s Alexa and Apple’s Siri, are susceptible to inaudible attacks [11, 12, 13]. Such attacks modulate voice commands onto ultrasonic carriers, generating commands imperceptible to human ears but can be accepted by voice control systems. Despite some levels of success in specific scenarios, existing portable inaudible attacking methods (such as relying on lithium batteries and using small speaker arrays to transmit attack signals) [13, 18, 12] have a short attack range of up to 3 m. However, a close-in attack within 3 m can be easily detected by many vigilant individuals [19]. Extending this range by simply increasing transmission power would lead to audible noise, thereby compromising the stealthiness of the attack. To address this limitation, more recent work, as represented by LipRead [19], divides the spectrum of inaudible commands to reduce audible leakage and employs a large array of 61 speakers to extend the attack distance up to 7.62 meters. DolphinAttack [20], targeting the iPhone X, uses 40 speakers along with external power supplies and amplifiers to achieve a range of 19.8 meters. In comparison, although these methods can significantly expand the attack distance, their reliance on bulky equipment and external power sources greatly reduces portability, limiting their applicability in many real-world public settings.
@@
Figure 1: A possible attacking scenario of MetaAttack where an attacker can launch an inaudible attack in a public place to remotely control the target device in another building. In this scenario, the attack device must be stealthy and does not require the use of other continuously powered devices such as DC power supply.
@@
We showcased that MetaAttack can be implemented with low-cost, commercial, off-the-shelf components like Raspberry Pi, amplifiers, loudspeakers, servo, and a 3D-printed MAM. We thoroughly evaluate MetaAttack under different practical scenarios under different attack ranges, tasks, noise levels, obstacles, and targeting different target devices. We compare MetaAttack against three prior inaudible attacks [12, 13, 19]. Experimental results show that MetaAttack is robust under different evaluation scenarios and tasks and can successfully launch an inaudible attack beyond 8 m range with a noise level of up to 65 dB. Compared with prior attacking methods, MetaAttack achieves a 76% average success rate under a 8.85 m attack range, making it the smallest device to achieve such performance.
@@ II-A Inaudible Attacks
In an inaudible attack, the microphone receives input signals that include not only the audible-frequency voice control signal but also a high-frequency carrier signal. Due to the microphone’s non-linear response, these input signals generate additional non-linear components in the output, providing the conditions for recognizing and transmitting the attack signal. Eq.1 represents the composition of the input signal ${S_{in}}(t)$, which is as follows:
@@
where ${V_{attack}}\left(t\right)$ is the voice control signal, typically a low-frequency control signal that the device can recognize. $\cos\left({2\pi{f_{carrier}}t}\right)$ is the high-frequency carrier signal, with a frequency of $f_{carrier}$}, which is beyond the human hearing range (typically above 20 kHz). The role of the carrier signal is to add additional frequency components to the attack signal, utilizing the non-linear characteristics of the speaker array when outputting the signal to generate additional components. The attack signal output by the speaker array, ${S_{attack}}(t)$, is represented by Eq.2:
@@
where $S_{h}$ represents background noise or other unrelated signals that are not associated with the attack. The term $\frac{A_{2}}{2}\left(1+V_{attack}^{2}(t)+2V_{attack}(t)\right)$ represents the additional signal components introduced by non-linear effects. In this case, ${V_{attack}}\left(t\right)$ can be recognized by the target’s microphone, enabling the successful execution of the inaudible attack.
@@ II-B Threat Model
Portable, miniaturized and stealthy inaudible attack. We focus on implementing an inaudible acoustic attack that is highly portable, miniaturized, and can be executed outside of detection range. Miniaturization and portability mean that the system does not rely on signal amplifiers powered by DC power supplies or bulky speaker arrays, and can be easily carried in a handbag or small backpack, allowing the attacker to operate discreetly in public spaces. The so-called detection range refers to a range in which the target is unlikely to detect the presence of the attacker. According to related studies [39, 40, 41], a distance greater than 3 meters is considered a safe distance when interacting with strangers, typically not raising suspicion.
No victim device interaction during attack. We assume that during the attack, the victim’s device is in close proximity but not actively in use, with its voice assistant idle and the screen not being viewed by the victim. This assumption aligns with those made in prior audio-based attack studies, such as Dolphin [12, 20], NUIT [42] and GhostTalk [15], and reflects realistic scenarios, such as when the victim places the device aside or face-down on a desk to focus on tasks like studying or attending a meeting.
Obtain the victim’s voiceprint. We assume the attacker can synthesize the target’s voice using, e.g., deep-learning-based voice generation from a few training samples [12]. We believe this assumption is reasonable in many day-to-day scenarios where the target’s voice samples can be collected from public places, a multi-user meeting, or audio and video clips available to the attacker.
Attack victim device in pocket. We consider the scenario where the victim may place their device in their pocket for convenience while out, during which the pocket could physically interfere with the transmission of attack commands. For this scenario, we have verified the system’s effectiveness in Section VI-C2.
@@ II-C Practical Limitations of Prior Methods
Launching inaudible voice commands is achieved by modulating the control signal, but high-power operation of speakers can cause nonlinear effects, leading to audible leakage, which limits the attack range. Previous studies [12, 13] avoided leakage by limiting the device power, but this restricted the attack range to within 3 meters. To achieve longer-range attacks, LipRead [19] divides the frequency spectrum into low-frequency parts to reduce human perception, but it requires 61 speakers, increasing the risk of detection, and spectrum segmentation may reduce the accuracy of commands. Similarly, Yan et al. [20] extended the attack range to 19.8 meters, but this requires a signal generator and DC power supply, which does not meet the needs of portable attacks, and requires 40 speakers. To clearly illustrate the limitations of existing research, we provide a comparison in Table I. Most studies focus on voiceprint recognition attacks [19, 11, 13, 12], but these attacks often require close proximity or a large number of auxiliary devices, lacking stealth. Therefore, a successful attack requires a balance between portability and attack range.
@@
System Setup Power supply Signal generator Speakers Attack range Portable Stealthy
DolphinAttack (Long-range System) [20] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $26*16.5*30~cm\approx$ 12MacBookPro)12~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $42.6*8.8*48.3~cm\approx$ 16MacBookPro)16~MacBook~Pro) 40 19.8m (iPhoneX) 10m (iPhoneSE) No No
LipRead [19] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $1.3*31.2*36~cm\approx 22$MacBookPro)22~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3~cm\approx$ 7MacBookPro)7~MacBook~Pro) 61 7.62m No No
BackDoor [11] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $1.3*31.2*36~cm\approx 22$MacBookPro)22~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3~cm\approx$ 7MacBookPro)7~MacBook~Pro) 18 3.5m No No
EchoAttack [13] [Uncaptioned image] No Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3cm\approx$ 7MacBookPro)7~MacBook~Pro) 40 2.23m Yes Yes
@@ III OVERVIEW of OUR APPROACH
Figure 3: Overview of MetaAttack with three steps.
@@
Processing inaudible commands. The initial phase of our attack involves processing and modulating voice commands. We first capture the victim’s voice and synthesize their voiceprint (see Section II-B). Then, we modulate the voice commands generated by Google Text-to-Speech AI [45] together with the synthesized voiceprint onto ultrasonic carrier waves using the method described in Section II, generating inaudible commands [19]. Each command group includes four types of inaudible instructions: Mute commands (e.g., “speak 6%” for Siri or “turn the volume to 1” for other devices [42, 46], denoted as $MuteCmd$), Attack commands ($AttackCmd$), Feedback commands (e.g., “turn on hotspot” and “turn off hotspot,” denoted as $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$) and Reset commands (e.g., “speak 50%” for Siri or “turn the volume to 5” for other devices, denoted as $ResetCmd$).
Attacking target device. To accurately target the device, MetaAttack uses a servo motor to rotate in 12° steps within a specified range, gradually aligning with the victim. The system first sends the $MuteCmd$ to mute the device [42, 46], then sends the $AttackCmd$ to carry out the attack, activates the feedback mechanism to confirm the success of the attack, and finally uses the $ResetCmd$ to restore the device’s volume. Leveraging recent research findings[47, 48, 49], our system has successfully implemented the covert deletion of call and text message records, thereby significantly enhancing the attack’s stealth capabilities.
To ensure the reliability of the attack execution, each command is sent consecutively five times. Since the execution time of each command is only a few seconds, repeated transmissions do not introduce any significant negative effects. Leveraging MetaAttack’s high single-command success rate of 76% at a distance of 8.85 meters, this strategy raises the theoretical final success rate of each command to 99.76%. In addition, we conducted 50 rounds of complete feedback processes to validate the execution success rate of each attack stage. Among them, the $MuteCmd$, AttackCmdAttackCmd, FeedbackCmdsFeedbackCmds and $ResetCmd$ each achieved a 98% success rate, with only one failure per command. The overall success rate of the full attack chain was 94%, with only three failures, demonstrating the high efficiency and stability of the attack process.
Activating feedback mechanism. This feedback mechanism uses a Raspberry Pi to monitor the state changes of the victim device after receiving a wireless switching command to determine whether the attack was successful. If the victim device’s hotspot signal undergoes a state transition of appearancedisappearancereappearance, the command execution is considered successful.
The reason for using the sequence appearancedisappearancereappearance is that relying solely on appearancedisappearance may lead to false positives. If a non-victim device happens to activate its hotspot between the two feedback commands, the system may incorrectly interpret this as an attack success and prematurely terminate the attack process. To address this, we introduce appearancedisappearancereappearance to confirm the continuity of the response, as it is highly unlikely for a non-victim device to switch its hotspot twice within a short period. To further reduce false positives, we incorporate iwlist [50] to monitor hotspot SSIDs and signal strength in real time, addressing the random appearance or disappearance of weak signals from non-victim devices. Typically, such weak signals show gradual changes in strength when appearing or disappearing, whereas the signal from the victim device triggered by $FeedbackCmd$ exhibits sudden changes. Based on this difference, we filter out SSIDs with abrupt signal strength transitions, effectively eliminating interference and improving detection accuracy.
@@
As shown in Algorithm 1 and Fig. 4, to illustrate the feedback mechanism for smartphones, we break it down into four sequential steps:
Step 1. After transmitting the $MuteCmd$ and $AttackCmd$ (line 1), the Raspberry Pi uses iwlist to record the identified hotspot devices and stores them in the L1 list (line 1). Then, it sends the $FeedbackCmd$ and stores the new records in the L2 list (line 1). To prevent the system from repeatedly executing $MuteCmd$ and $AttackCmd$ while waiting for the success of $FeedbackCmd$, we introduced the concept of a time window. Considering that the system needs a certain amount of time to receive the victim device’s hotspot signal via the Raspberry Pi, no other commands will be sent within a time window after sending $FeedbackCmd$, thereby avoiding the repeated execution of commands.)
@@
Step 2. The system continues to send $FeedbackCmd_{2}$ and stores the received hotspot information in list L3 (line 1). If there are discrepancies in the hotspot signals ($dif_{1}$) between the L2 and L3 lists, the process moves to Step 3 (line 1). Otherwise, the MetaAttack is rotated to the next angle (line 1).
@@
Step 3. MetaAttack will sequentially send $FeedbackCmd_{1}$ and then instruct the Raspberry Pi to retrieve the L4 hotspot list (line  1). If there are differences in the hotspot signals ($dif_{2}$) between the L3 and L4 lists, a successful stealth attack is detected (lines 1 - 1). If not, the MetaAttack will be adjusted to the next angle.
@@
Step 4. To maintain the stealth of the attack, this step focuses on restoring the victim device to its original state before the attack (line 1).
Optimization of feedback strategies in different scenarios. Considering the time overhead of the feedback algorithm and the limited angular coverage of a single attack, we have designed additional strategies to improve attack efficiency: for attacks on visible targets, the attacker can manually align the system with the target, completing the attack within 1 minute.
In addition, for commands that provide direct feedback to the attacker (such as making a call for eavesdropping), we immediately confirm whether the call was successfully connected after the command is executed. If the connection is successful, the attack is stopped via SSH connection to the Raspberry Pi; if it fails, the subsequent feedback is halted and the system moves to the next angle to reduce time overhead. By tailoring strategies to different environments, the system achieves efficient and flexible attack execution.
@@
To confirm the feasibility of our attack process, we have released a complete process demonstration [36] and open-sourced the feedback code [38].
Figure 4: MetaAttack Overview: MuteCmdMuteCmd and AttackCmdAttackCmd initiate the attack, FeedbackCmdFeedbackCmd confirms success and ResetCmdResetCmd restores the target.
1
2foreach angleangle do
3
4 SendMuteCmd,AttackCmd,activateiwlistSend\ MuteCmd,\ AttackCmd,\ activate\ iwlist
5